OpenStudy (blaque678):

Determine the average rate of change of the function f(x)=x+x^3 between x = 0 and x = 3

zepdrix (zepdrix):

So the average rate of change will be given by:\[\Large \frac{f(3)-f(0)}{3-0}\]

zepdrix (zepdrix):

\[\Large f(\color{royalblue}{x})\quad=\quad (\color{royalblue}{x})+(\color{royalblue}{x})^3\] Do you understand how we find f(3) and f(0)? Maybe the colors will help.\[\Large f(\color{royalblue}{3})\quad=\quad (\color{royalblue}{3})+(\color{royalblue}{3})^3\]\[\Large f(\color{royalblue}{0})\quad=\quad (\color{royalblue}{0})+(\color{royalblue}{0})^3\]

OpenStudy (blaque678):

30 and 0

zepdrix (zepdrix):

Mmm yah that sounds right :) So what do you get when you plug all of the pieces in?

OpenStudy (blaque678):

0 and 30 when I plugged it in

zepdrix (zepdrix):

Ya 30 and 0 sound right :) Plug them in like this and simplify. \[\Large \frac{f(3)-f(0)}{3-0}\qquad=\quad \frac{30-0}{3-0}\]